OpenStudy (anonymous):

the avg age of 8 men is decreased by 2 years when two of them whose ages are 22 and 28 are replaced by two boys, what is the avg age of the two boys?

OpenStudy (anonymous):

@wasiqss Sorry computer hanged

OpenStudy (wasiqss):

np , chalo isse bi krte hain

OpenStudy (anonymous):

17

OpenStudy (anonymous):

Sheg answer mujhe bhi pata hai, peche answer key me dia hua hai.

OpenStudy (anonymous):

but book tumhare pas hai mere pas nahi

OpenStudy (wasiqss):

@Zeerak dimag thanda rakhiye janab :P

OpenStudy (anonymous):

see total age kitna kam hua 8*2 = 16 yrs age of replaced boys = 22 + 28 = 50 Age of new boys = 50 - 16 = 34 Avg of New Boys = 34/2 = 17

OpenStudy (wasiqss):

zeerak samajh aya?

OpenStudy (anonymous):

Thora time do

OpenStudy (wasiqss):

ok samjho